Write (9/4)^-1 as an exponent

Answer by nabla(475)

If we take something to the negative power, it is the same as the reciprocal to the same, but positive, power. So,
(9/4)^-1 =(4/9)^1=4/9
You could write this as (2/3)^2 as well.